Problem 8-06 You expect to invest your funds equally in four stocks with the following expected returns: Stock Expected Return A 15 % B 13 C 10 D 5 At the end of the year, each stock had the following realized returns: Stock Realized Return A -2 % B 29 C 2 D -2 Compare the portfolio’s expected and realized returns. Round your answers to two decimal places. Expected return: %? Realized return: %?
Answer #1

Expected/realized return from portfolio is the weighted average return

Therefore,

Expected return from portfolio = (15+13+10+5)/4 =10.75%

Realized return from portfolio = (-2+29+2-2)/4 = 6.75%